OpenStudy (anonymous):

please help If x - 7 = 10, then x - 7 + 7 = 10 + 7. This is an example of which property below?

OpenStudy (lgbasallote):

look closely: x - 7 = 10 then x - 7 + 7 = 10 + 7 what do you think changed?

OpenStudy (anonymous):

i know what changed i just dont know what property it is

OpenStudy (lgbasallote):

if you know what changed...then you'll know what property it is....so if you don't know what property it is...that means you don't know what changed so what changed?

OpenStudy (anonymous):

is it the Commutative Property for Addition

OpenStudy (lgbasallote):

do you know what changed or not?

OpenStudy (anonymous):

yeah the 7 you add it to both sides

OpenStudy (lgbasallote):

that's right. so if you add to both sides...it's called addition property

OpenStudy (lgbasallote):

commutative property is this a + b = b + a

OpenStudy (lgbasallote):

commutative property is changing the order

OpenStudy (anonymous):

oh okay i see thank you (:

OpenStudy (lgbasallote):

welcome